Convert 150 lbm to a) micrograms and b) tonnes


1
Expert's answer
2021-03-31T09:39:57-0400

a)


"150\ lb_m=150(0.4536)(10^9)=6.8\cdot10^{10} \mu g"

b)


"150\ lb_m=150(0.4536)(10^{-3})=0.068\ tonnes"